The odds against an event are 8:8. Find the probability that the event will occur.
Found 3 solutions by ewatrrr, sudhanshu_kmr, solver91311:
Answer by ewatrrr(24785)


Hi
The odds against an event are 8:8
In this case then, odds for would be 8:8
P(that the event will occur) = 8/16 = 1/2

Answer by sudhanshu_kmr(1152)


odds against the event are 8:8
probability = 8/(8+8) = 8/16

= 1/2

50% probability........



Answer by solver91311(24713)



The odds against an event means that the event will not occur times for every times that it does occur.

Hence the probability that the event will occur is:



And the probability that the event will NOT occur is:



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism